# bibliography.bib

### v2

@@ 4.4 Adaptive MPML algorithm
In this section we develop an adaptive algorithm which will automatically choose a suitable effective precision on each level of the MLMC method. We will use the standard MLMC algorithm as the foundation for our proposed algorithm. The first step is the choice of the target effective precision.
In order to choose the correct effective precision in each step, we will use the error bound for the MPML method from Theorem 4.1. For simplicity, we assume the bias and variance decay as in (14). We propose the following approach: choose the accuracy $\delta_{l}$ on level $l$ such that the total MSE of the MPML estimator is not greater than a constant times the MSE of the standard MLMC estimator for a fixed constant $k_{p}\in(0,1)$, the choice of which we discuss later. According to Theorem 4.1, for this to hold it suffices to choose $\delta_{l},\;l=0,\ldots,L$ such that
@@
With this choice, both inequalities in (23) are satisfied and we obtain the desired error estimate if the rest of the assumptions in Theorem 4.1 are satisfied. The reason we “hide” the computational error rather than optimally balancing it with the model error in our adaptive algorithm is that in the case when the computational error comes from a linear solver, it typically decays exponentially with the number of iterations (e.g., in iterative refinement; see Section 2.3) and therefore balancing the two errors would not bring great benefits. However, it might be of interest in cases when the computational error decreases polynomially.
@@
Let us discuss in more detail the choice of the constant $k_{p}$. Although in this work the constant $k_{p}$ is chosen to be fixed, more general choices are possible, one of which is discussed in the next paragraph. Note that the accuracy $\delta_{l}$ defined by (24) is, for $l<L$, not sensitive with respect to the changes of the constant $k_{p}$. Indeed, if the constant is decreased 100×100$\times$100 ×, the target accuracy decreases only 10×10$\times$10 × for our model problem; see Theorem 4.4. The value $k_{p}\coloneqq 0.05$ is a safe choice to bound the computational error, as demonstrated in Section 5. Note also that the values of $\delta_{l}$ can be computed “on the fly” with no additional cost, given that the decay rates of bias and variance in Theorem 4.1 are known. We now proceed to the formulation of the MPML algorithm, given in Algorithm 4.1.
It is natural to ask how the choice of the constant $k_{p}$ affects the number of samples $N_{l}$ required on each level $l$ to achieve the desired tolerance. According to Theorem 4.1, if the accuracy $\delta_{l}$ is chosen according to (24) then the variance is increased on each level at most by approximately the factor $(1+k_{p})$. This means that the number of samples $N_{l}$ on each level is increased at most by the same factor $(1+k_{p})$; see step 7 of Algorithm 4.1 Since $k_{p}\ll 1$, this does not pose a problem for us. Depending on the exact settings of the problem (on the linear solver), it might make sense to use a different cost model for MPML than for MLMC to estimate the cost per sample $C_{l}$ on each level, which may impact the number of samples on each level as well. However, in Figure 3 we verify numerically that in our example the overall increase in the number of samples compared to standard MLMC is negligible.
Throughout this section we have assumed that we are able to compute the approximate discrete solution $\widehat{u}_{h_{l}}$ on the level $l$ effectively to precision $\delta_{l}$. How to achieve this, depends on the inexact solver that is used for the linear system. We discuss here two basic cases, iterative and direct solvers. First, any suitable iterative solver can in principle be used with the stopping criterion given by (13). Since the values of $\delta_{l}$ obtained using the adaptive MPML algorithm are typically relatively “big” (see Section 5 for examples), the iterative solver can potentially achieve the tolerance in a very small number of iterations, leading to significant cost gains. In this case, the cost gain does not come primarily from using low precision, but rather from reducing the number of iterations. As an example, we employ MINRES; see Section 2.2.
The second technique we propose to use, is iterative refinement. It is used here in combination with a direct solver based on Cholesky factorisation. Iterative refinement can, however, be used in principle with any direct or iterative solver; see Section 2.3 and [carson2018accelerating]. As explained in Section 2, iterative refinement helps us to achieve just the desired accuracy with a small additional cost.
@@ 4.5 Cost analysis
The cost gain using the adaptive MPML algorithm from Section 4.4 depends on which of the three complexity regimes of the MLMC estimator in [cliffe2011multilevel, Theorem 1] applies. These depend on the relative sizes of $\beta$ and $\gamma$ in (10).
@@
The abstract cost analysis is done here in terms of arbitrary cost units. In applications, we may consider, e.g., CPU time, memory references, or floating point operations, depending on what best fits our purpose. At the end of this subsection we apply the abstract cost analysis to the Problem 4.2 and give a concrete example of cost measures for this problem.
For the purpose of the abstract analysis we assume the following. For the standard MLMC we assume $\operatorname{\textrm{var}}[Y_{l}]=c_{v}m^{-\beta l}$ and $C_{l}=c_{c}m^{\gamma l}$ with $\beta>\gamma$; see (10) and [cliffe2011multilevel, Theorem 1]. Further, we assume that both MLMC and MPML algorithms use the same number of samples $N_{l}$ on each level and the variance on each level $\operatorname{\textrm{var}}[Y_{l}]$ is the same for both algorithms. This is a reasonable assumption, since our adaptive MPML algorithm chooses the computational error significantly smaller than the model error (see the discussion in Section 4.4 and Figure 3). For the costs per sample on each level with low accuracy we assume only that
@@
As an example, consider Problem 4.2 with uniform mesh refinement by a factor of $m=2$. Using an optimal linear solver (e.g., multigrid), we have $\gamma=2$ and the variance of $Y_{l}$ decays with a rate of $\beta=4$ (see [cliffe2011multilevel]). If $q=1/4$, for example, then Corollary 4.5 predicts that the cost is reduced by at least a factor of 1.6. In our experiments, using the MINRES method, we observed an actual cost gain in terms of floating point operations by a factor of $\approx 1.5$ (see Figure 2). When a low precision sparse direct solver with iterative refinement is used, the reduction in allocated memory is reduced by up to a factor of $\approx 3.5$ (see Figure 8).
@@ 5 Numerical results
We provide numerical experiments demonstrating the potential cost savings of the adaptive MPML algorithm (Algorithm 4.1) over a standard adaptive MLMC algorithm. The adaptive MPML algorithm preserves the overall computational accuracy. As a suitable model, we use an elliptic PDE with lognormal random coefficients with both a direct and an iterative linear solver. All numerical experiments are implemented in Python. The codes are available at https://github.com/josef-martinek/mpml.
@@ 5.1 Elliptic PDE with lognormal coefficients – iterative solver
Here $\omega=(\omega_{1},\ldots,\omega_{s})\in\mathbb{R}^{s}$ is such that $\omega_{j}\sim N(0,\sigma^{2})$ for a fixed $\sigma>0$. Random fields of this form are widely used; see [babuska2004galerkin], [nobile2008sparse], and [cliffe2011multilevel] for examples.
@@
00 $hhss$
$1$ $ssss$
$2+$ $ssdd$
@@ 00 $hhss$
Table 1: Left: Required effective precision on each level (in terms of relative residual), determined by the adaptive MPML algorithm (Algorithm 4.1) for different values of the finest level $L$.  Right: Choice of precision values $(\varepsilon_{f},\varepsilon_{7},\varepsilon,\varepsilon_{r})$ for iterative refinement on each level $l$. The factorisation is carried out in half precision on level 00.
As the quantity of interest we choose
Due to the fact that $\omega_{j}\sim N(0,\sigma^{2})$, assumptions (4) and (5) are not satisfied. By choosing, e.g., $\omega_{j}\sim\text{Uni}(0,c)$, one could ensure that both assumptions are satisfied, but we want to test the developed methods in a more realistic setting. The analysis in Section 3.2 could easily be extended to this setting, using the analysis for MLMC presented in [teckentrup2013further], but we did not present this here to avoid unnecessary technicalities.
@@
In the first example of this section, we choose the data in (27) as follows. The right-hand side satisfies $f\equiv 1$ on $D$ and the parameters in the coefficient function are chosen as $s=4$, $q=2$, and $\sigma=2$. To solve this problem numerically, we discretise (27) using the FEM as described in Section 3 with simplical elements implemented in FEniCSx [baratta_2023_10447666]. In this experiment, the discretisation parameter on the coarsest mesh is $h_{0}=1/8$ and the mesh is refined on each level by a factor $m=2$.
@@
Recall that in all our experiments we refer to standard MLMC as Algorithm 4.1 without step 2, where the computations in step 4 are carried out with an a-priori given, sufficiently high accuracy. In each experiment we always describe precisely what this accuracy is. For MLMC, the parameters $\alpha$ and $\beta$ in the MLMC complexity theorem [cliffe2011multilevel, Theorem 1] (see also (10)) can be chosen to be $\beta=4$, $\alpha=2$ for the random field in (28). To set up the MPML algorithm (Algorithm 4.1), we further need the parameters $\alpha_{1}$, $\alpha_{2}$, $\beta_{1}$, and $\beta_{2}$ from the MPML complexity theorem (Theorem 4.1). As shown in Lemma 4.3, we have $\alpha_{1}=\alpha=2$, $\beta_{1}=\beta=4$, $\alpha_{2}=1$, and $\beta_{2}=2$. In both adaptive algorithms, the underlying linear systems are solved using the PETSc implementation of MINRES [petsc_4_py] with a stopping criterion given by the relative residual norm. The tolerances for the stopping criterion in the MPML algorithm are given in Table 1 (left). The standard MLMC algorithm uses a fixed tolerance specified below for each experiment and all computations are carried out in double precision without iterative refinement. The cost of solving the linear system is estimated in terms of the number of floating point operations (FLOPs) performed by MINRES. To count the FLOPs we use the PETSc GetFlops() function.
@@
At the beginning of the cost analysis in Section 4.5 we made the assumption that the numbers of samples on each level are chosen approximately the same within the MPML and the MLMC adaptive procedures. We verify numerically that this is true. Figure 3 shows the average number of samples obtained by MPML (Algorithm 4.1) and standard MLMC on each level for the tolerance $\text{TOL}^{2}=10^{-6}$. We observe that the number of samples are equal within statistical error given by the approximate $95\%$ confidence intervals. For this experiment, the standard MLMC algorithm uses a fixed tolerance of $10^{-10}$ in the MINRES stopping criterion. Note that Level 4 is not reached in all runs and we need on average only one sample on Level 3.
In Section 4.4 we claimed that the MPML algorithm is not sensitive to the a-priori chosen constant $k_{p}$ used to determine the required computational accuracy. Figure 4 shows the average MSE of the MPML algorithm (Algorithm 4.1) for the MSE tolerance $\text{TOL}^{2}=2\times 10^{-6}$ with approximate $95\%$ confidence intervals. The MSE is estimated using $1000$ runs of the algorithm for the values $k_{p}=0.05,0.1,0.2,0.4$. The estimated MSE of the standard MLMC algorithm is shown for comparison with a dashed line. It is estimated using $1000$ runs of standard MLMC with a tolerance of $10^{-10}$ in the MINRES stopping criterion. The choice of the constant $k_{p}$ seems to have no significant impact on the overall accuracy, provided the constant is chosen sufficiently small. For our suggested choice of $k_{p}=0.05$ the computational error in our model problem is bounded safely.

### v3

@@ 4.3 Adaptive MPML algorithm
In this section we develop an adaptive algorithm which will automatically choose a suitable computational accuracy on each level of the MLMC method. We will use the standard MLMC algorithm as the foundation for our proposed algorithm.
In order to choose the correct computational accuracy in each step, we will use the error bound for the MPML method from Corollary 4.2. We propose the following approach: choose the accuracy $\delta_{l}$ on level $l$ such that the total MSE of the MPML estimator is not greater than a constant times the MSE of the standard MLMC estimator for a fixed constant $k_{p}\in(0,1)$, the choice of which we discuss later. According to Corollary 4.2, for this to hold it suffices to choose $\delta_{l},\;l=0,\ldots,L$ such that
@@
With this choice, both inequalities in (15) are satisfied and we obtain the desired error estimate from Corollary 4.2.
Remark 4.4 (Balancing the computational and model error).
The reason we “hide” the computational error rather than optimally balancing it with the model error in our adaptive algorithm is that in the case when the computational error comes from a linear solver, it typically decays exponentially with the number of iterations (e.g., in iterative refinement; see Section 2.3) and therefore balancing the two errors would not bring great benefits. However, it might be of interest in cases when the computational error decreases polynomially.
@@
Let us discuss in more detail the choice of the constant $k_{p}$. Although in this work the constant $k_{p}$ is chosen to be fixed, more general choices are possible in principle. The value $k_{p}\coloneqq 0.05$ is a safe choice to bound the computational error, as demonstrated in Section 5. Note also that the values of $\delta_{l}$ can be computed “on the fly” with no additional cost, given that the decay rates of bias and variance in (14) are known. For the resulting adaptive MPML algorithm see Algorithm 4.1.
It is natural to ask how the choice of the constant $k_{p}$ affects the number of samples $N_{l}$ required on each level $l$ to achieve the desired tolerance. According to Corollary 4.2, if the accuracy $\delta_{l}$ is chosen according to (16) then the variance is increased on each level at most by approximately the factor $(1+k_{p})$. This means that the number of samples $N_{l}$ on each level is increased at most by the same factor $(1+k_{p})$; see step 7 of Algorithm 4.1. Since $k_{p}\ll 1$, this does not pose a problem for us. Depending on the exact settings of the problem (on the linear solver), it might make sense to use a different cost model for MPML than for MLMC to estimate the cost per sample $C_{l}$ on each level, which may impact the number of samples on each level as well. However, in Figure 4 we verify numerically that in our example the overall increase in the number of samples compared to standard MLMC is negligible.
@@ 4.4 Cost analysis
The cost gain using the adaptive MPML algorithm from Section 4.3 depends on which of the three complexity regimes of the MLMC estimator in [8, Theorem 1] applies. These depend on the relative sizes of $\beta$ and $\gamma$ in (11).
@@
The abstract cost analysis is done here in terms of arbitrary cost units. In applications, we may consider, e.g., CPU time, memory references, or floating point operations, depending on what best fits our purpose. At the end of this subsection we apply the abstract cost analysis to the Problem 4.6 and give a concrete example of cost measures for this problem.
For the purpose of the abstract analysis we assume the following. For the standard MLMC we assume $\operatorname{\textrm{var}}[Y_{l}]=c_{v}m^{-\beta l}$ and $C_{l}=c_{c}m^{\gamma l}$ with $\beta>\gamma$; see (11) and [8, Theorem 1]. Further, we assume that both MLMC and MPML algorithms use the same number of samples $N_{l}$ on each level and the variance on each level $\operatorname{\textrm{var}}[Y_{l}]$ is the same for both algorithms. This is a reasonable assumption, since our adaptive MPML algorithm chooses the computational error significantly smaller than the model error (see the discussion in Section 4.3 and Figure 4). For the costs per sample on each level with low accuracy we assume only that
@@
4.5 Application to the elliptic PDE problem
Problem 4.6.
Under assumptions (4) and (5), it can be shown that when MLMC is applied to Problem 4.6, we obtain (11) with $\alpha=2$ and $\beta=4$. Generally $\gamma$ depends on the linear solver used and we discuss it in our numerical experiments. If an optimal linear solver is used (e.g., multigrid), one has $\gamma=2$; see [8].
It can be shown that the MPML bias and variance decay assumptions (14) are satisfied for Problem 4.6. The corresponding values of the constants $\alpha_{1}$, $\alpha_{2}$, $\beta_{1}$, and $\beta_{2}$ are given by the following lemma.
Lemma 4.7.
$|\mathbb{E}[\tilde{Q}_{l}-Q]|$ $=O(h_{l}^{2}+\delta_{l}),$ (20)
$\operatorname{\textrm{var}}[\tilde{Y}_{l}]$ $=O(h_{l}^{4}+\delta_{l}^{2}).$ (21)
Proof.
Using Jensen’s inequality the bias error in (20) can be decomposed as
From [8, Section 3] it follows that
$\|G(u)-G(u_{h})\|_{L^{1}(\Omega,\mathbb{R})}\leq Ch^{2},$ (23)
The variance bound in (21) can be shown similarly. Let us estimate
$\leq\mathbb{E}\bigl[\bigl(\tilde{Q}_{l}-Q_{l}+Q_{l}-Q+Q-Q_{{l-1}}+Q_{{l-1}}-\tilde{Q}_{l-1}\bigr)^{2}\bigr].$
Using the same technique as in [8], we obtain an estimate of the form
$\operatorname{\textrm{var}}[\tilde{Y}_{l}]\leq$ $C\bigl(\mathbb{E}[(\tilde{Q}_{l}-Q_{l})^{2}]+\mathbb{E}[(Q_{l}-Q)^{2}]$
$+\mathbb{E}[(Q-Q_{{l-1}})^{2}]+\mathbb{E}[(Q_{{l-1}}-\tilde{Q}_{l-1})^{2}]\bigr).$ (25)
As in [8], the quantity $\mathbb{E}[(Q_{l}-Q)^{2}]+\mathbb{E}[(Q-Q_{{l-1}})^{2}]$ can be estimated by
$\mathbb{E}[(Q_{l}-Q)^{2}]+\mathbb{E}[(Q-Q_{{l-1}})^{2}]\leq Ch_{l}^{4}.$ (26)
This lemma allows us to formulate a specific version of Corollary 4.2 regarding the error of the MPML estimator applied to Problem 4.6 discretised using finite elements. We summarize this in the following corollary.
Corollary 4.8 (Error of the MPML FEM).
Proof.
The claim follows from Lemma 4.7 and Corollary 4.2. ∎
Lemma 4.7 now allows us to specify the accuracy parameters $\delta_{l}$ of the adaptive MPML algorithm from Section 4.3 for Problem 4.6. Using Lemma 4.7 in (16) we get
$\begin{split}\delta_{l}&\coloneqq\sqrt{k_{p}}h_{l}^{2},\quad l=0,\ldots,L-1,\\ \delta_{L}&\coloneqq k_{p}h_{L}^{2}.\end{split}$ (28)
To quantify the cost gains, let us apply the abstract analysis from Section 4.4 to Problem 4.6. Using an optimal linear solver (e.g., multigrid), we have $\operatorname{\textrm{var}}[Y_{l}]=O(2^{-4l})$ and $C_{l}=O(2^{2l})$ (see [8]). If in (17) we have $q=1/4$, for example, then Corollary 4.5 predicts that the cost is reduced by at least a factor of 1.6.
In our experiments, using the MINRES method, we observed an actual cost gain in terms of floating point operations by a factor of $\approx 1.5$ (see Figure 3). When a low precision sparse direct solver with iterative refinement is used, the reduction in allocated memory is reduced by up to a factor of $\approx 3.5$ (see Figure 9).
Remark 4.9 (Application to other sampling methods).
Note that the extension of the standard MLMC assumptions (11) to the more general MPML error model (12) and (13) is not specific to multilevel Monte Carlo. Also, Lemma 4.7 is not specific to the sampling method. It is a statement about the quantity of interest.
In fact, assumptions of the form (11) are central to the analysis of many other multilevel sampling methods, such as multilevel MCMC [11]. The assumptions [11, Theorem 3.4, Assumptions M1., M2., M3.] for the analysis of multilevel MCMC are analogous to (11) and the method is also tested on a similar model problem; see [11, eq. (4.2)]. They can be extended in the same way as (11) to (20) and (21) to obtain a performance gain in multilevel MCMC via inexact computations and mixed precision arithmetic. As a second example, consider multi-index Monte Carlo [18]. An analogy to (11) for MIMC is [18, Assumptions 1 to 3]. An extension of these assumptions to take into account computational error is again straightforward.
@@ 5 Numerical results
We provide numerical experiments demonstrating the potential cost savings of the adaptive MPML algorithm (Algorithm 4.1) over a standard adaptive MLMC algorithm. The adaptive MPML algorithm preserves the overall computational accuracy. As a suitable model, we use an elliptic PDE with lognormal random coefficients with both a direct and an iterative linear solver.
In principle, any suitable iterative solver with the stopping criterion given by (19) can be used to compute the approximate discrete solution $\widehat{u}_{h_{l}}$ on the level $l$ effectively to precision $\delta_{l}$. Since the values of $\delta_{l}$ obtained using the adaptive MPML algorithm are typically relatively “big” (see Section 5.1 for examples), the iterative solver can potentially achieve the tolerance in a very small number of iterations, leading to significant cost gains. As an example, we employ MINRES as described in Section 2.2. In this case, the cost gains do not come primarily from using low precision, but rather from reducing the number of iterations. A technique, where the cost gains come from the use of low precision, is iterative refinement. It can in principle be used with any direct or iterative solver; see Section 2.3 and [6]. Here, we used it in combination with a direct solver based on Cholesky factorisation.
All numerical experiments are implemented in Python. The codes are available at https://github.com/josef-martinek/mpml.
@@ 5.1 Elliptic PDE with lognormal coefficients – iterative solver
Here $\omega=(\omega_{1},\ldots,\omega_{s})\in\mathbb{R}^{s}$ is such that $\omega_{j}\sim N(0,\sigma^{2})$ for a fixed $\sigma>0$. Random fields of this form are widely used; see [1], [26], and [8] for examples.
As the quantity of interest we choose
Due to the fact that $\omega_{j}\sim N(0,\sigma^{2})$, assumptions (4) and (5) are not satisfied. By choosing, e.g., $\omega_{j}\sim\text{Uni}(0,c)$, one could ensure that both assumptions are satisfied, but we want to test the developed methods in a more realistic setting. The analysis in Section 3.2 could easily be extended to this setting, using the analysis for MLMC presented in [30], but we did not present this here to avoid unnecessary technicalities.
@@
0 hhsshhss
11 ssssssss
$2+$ ssddssdd
@@
Table 1: Left: Required effective precision on each level (in terms of relative residual), determined by the adaptive MPML algorithm (Algorithm 4.1) for different values of the finest level $L$. Right: Choice of precision values $(\varepsilon_{f},\varepsilon_{7},\varepsilon,\varepsilon_{r})$ for iterative refinement on each level $l$. The factorisation is carried out in half precision on level 0.
@@
In the first example of this section, we choose the data in (29) as follows. The right-hand side satisfies $f\equiv 1$ on $D=(0,1)^{2}$ and the parameters in the coefficient function are chosen as $s=4$, $q=2$, and $\sigma=2$. To solve this problem numerically, we discretise (29) using the FEM as described in Section 3 with simplical elements implemented in FEniCSx [2]. In this experiment, the discretisation parameter on the coarsest mesh is $h_{0}=1/8$ and the mesh is refined on each level by a factor $m=2$.
@@
Recall that in all our experiments we refer to standard MLMC as Algorithm 4.1 without step 2, where the computations in step 4 are carried out with an a-priori given, sufficiently high accuracy. In each experiment we always describe precisely what this accuracy is. For MLMC, the parameters $\alpha$ and $\beta$ in the MLMC complexity theorem [8, Theorem 1] (see also (11)) can be chosen to be $\beta=4$, $\alpha=2$ for the random field in (30). To set up the MPML algorithm (Algorithm 4.1), we further need the parameters $\alpha_{1}$, $\alpha_{2}$, $\beta_{1}$, and $\beta_{2}$ from the MPML complexity theorem (Corollary 4.2). As shown in Lemma 4.7, we have $\alpha_{1}=\alpha=2$, $\beta_{1}=\beta=4$, $\alpha_{2}=1$, and $\beta_{2}=2$, which results in the effective precision choice given by (28). In both adaptive algorithms, the underlying linear systems are solved using the PETSc implementation of MINRES [9] with a stopping criterion given by the relative residual norm. The tolerances for the stopping criterion in the MPML algorithm are given in Table 1 (left). The standard MLMC algorithm uses a fixed tolerance specified below for each experiment and all computations are carried out in double precision without iterative refinement. The cost of solving the linear system is estimated in terms of the number of floating point operations (FLOPs) performed by MINRES. To count the FLOPs we use the PETSc GetFlops() function.
Figure 1: Computational error in the variance of the difference estimators $\operatorname{\textrm{var}}[\tilde{Y}_{l}]$ (left) and in the bias $|\mathbb{E}[\tilde{Q}_{l}-Q]|$ (right), plotted against effective precision. Discretisation level $l$ corresponds to $h_{l}=1/8\times 2^{-l}$. To achieve an effective precision $\delta_{l}$, relative residuals produced by MINRES are monitored.
We begin by visualising computational error in the variance $\operatorname{\textrm{var}}[\tilde{Y}_{l}]$ and in the bias $|\mathbb{E}[\tilde{Q}_{l}-Q]|$ (Figure 1). We plot both variance and bias against effective precision $\delta_{l}$. To achieve a given value of the effective precision $\delta_{l}$, we monitor the relative residuals produced by the MINRES method. A bound for the computational error in these quantities has been given in Lemma 4.7.
In Figure 1, we observe that the discretisation error quickly begins to dominate the variance and the bias on each level, as the computational error decays. On the fine levels, the variance decay rate (left) and the bias decay rate (right) become apparent. The initial decay rate is at least $O(\delta_{l}^{2})$ for the variance and at least $O(\delta_{l})$ for the bias, which are the rates obtained in Lemma 4.7. As the discretisation error starts to dominate, i.e., for small values of $\delta_{l}$ and large values of $h_{l}$, standard multilevel variance and bias are recovered. The sample variances have been produced using $10^{2}$ samples each, while for the bias estimates $10^{4}$ samples were used.
Let us note that the results in Figure 1 are not specific to the multilevel Monte Carlo method, only to the QOI. This shows the potential for using inexact computations in a broad spectrum of multilevel sampling methods if the QOI is sufficiently well-behaved, as discussed in Remark 4.9.
@@
At the beginning of the cost analysis in Section 4.4 we made the assumption that the numbers of samples on each level are chosen approximately the same within the MPML and the MLMC adaptive procedures. We verify numerically that this is true. Figure 4 shows the average number of samples obtained by MPML (Algorithm 4.1) and standard MLMC on each level for the tolerance $\text{TOL}^{2}=10^{-6}$. We observe that the number of samples are equal within statistical error given by the approximate $95\%$ confidence intervals. For this experiment, the standard MLMC algorithm uses a fixed tolerance of $10^{-10}$ in the MINRES stopping criterion. Note that due to the fact that the finest level is not specified a-priori in Algorithm 4.1 and due to the stochastic nature of the discrete choices in Algorithm 4.1, Level 4 is only reached in about 2 % of all runs. This suggests that in this example Level 4 might not be needed in practice. Note also that we need on average only one sample on Level 3.
In Section 4.3 we claimed that the MPML algorithm is not sensitive to the a-priori chosen constant $k_{p}$ used to determine the required computational accuracy. Figure 5 shows the average MSE of the MPML algorithm (Algorithm 4.1) for the MSE tolerance $\text{TOL}^{2}=2\times 10^{-6}$ with approximate $95\%$ confidence intervals. The MSE is estimated using $1000$ runs of the algorithm for the values $k_{p}=0.05,0.1,0.2,0.4$. The estimated MSE of the standard MLMC algorithm is shown for comparison with a dashed line. It is estimated using $1000$ runs of standard MLMC with a tolerance of $10^{-10}$ in the MINRES stopping criterion. The choice of the constant $k_{p}$ seems to have no significant impact on the overall accuracy, provided it is sufficiently small. For our suggested choice of $k_{p}=0.05$ the computational error in our model problem is bounded safely.